OpenStudy (vanz):

How do you solve this? x/5=x+6/x-2

OpenStudy (anonymous):

Yes these can definitively be tricky. Almost always do these first two steps 1) kill the fractions!!! 2) isolate x on one side Usually it will take you a couple of trys. x/5 = x+6/x-2 x = 5x+30 / x-2 multiply by 5 on each side x(x-2) = 5x +30 multiply by x-2 on each side x^2-2x = 5x+30 simplify x^2-2x-5x = 30 subtract 5x on each side x^2-7x = 30 simplify x^2-7x-30=0 subtract 30 on each side (x-10)*(x+3) = 0 factor x =10 and x =-3 here i factored. Im not sure if you know how to or not. If you dont then at the part where i factored use the quadratic formula instead